Time Frequency Brown Bag Seminar

Wednesday, April 29, 1998

12:00 noon

EQuad E415

Speaker: Frans M. Coetzee

Title: FILTERING OF OXIMETRY SIGNALS

Abstract:

I will discuss adaptive linear and non-linear filtering of physiological signals. The talk uses as a concrete example the design of an artifact resistant pulse oximeter. The principles of pulse oximetry and the characterization of the noise and signal components of patient data are briefly reviewed. An appropriate geometric framework for noise reduction which incorporates the physiological origin of the desired signal is formulated. Two approaches to solving the geometric problem are discussed. The first uses linear adaptive filtering where the signals are projected into an appropriately constructed subspace. The second approach discusses current work to perform nonlinear waveform fitting using globally convergent homotopy methods previously developed for global optimization of neural networks.